# thanks: 0000-0002-5209-872Xthanks: 0000-0003-2028-6782

### v1

@@ I Introduction
To address this issue, for long-baseline accelerator and reactor neutrino oscillations, we developed the NuFast-LBL algorithm [18] and the associated code on github [19]. This algorithm was based on a variety of theoretical neutrino oscillation results obtained over a number of years including some of the earliest full expressions for neutrino oscillations in constant matter [20, 21], solutions to the cubic equation [22, 18], the role of CP violation [23, 24, 25, 26, 27, 28], the optimal means of computing the eigenvectors [29, 30, 31, 32, 33, 18, 34], and the optimal structure for the oscillation probabilities as well as various other theoretical advances [35, 36, 37, 38, 39], see also [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. In this paper, we use parts of the NuFast-LBL algorithm and modify it to handle the varying density profiles through the Earth. The additional complexities of traversing sharply varying density profiles require a new approach constructed from the ground up, which is the result presented in this paper. We provide a new code called NuFast-Earth, also available on github, see [53] [^1] The code is available at github.com/PeterDenton/NuFast-Earth., which is significantly more powerful than the long-baseline (LBL) code. Notably, it intelligently saves reused aspects of the calculations for future uses which vastly speeds up many calculations of the oscillation probabilities in similar conditions. It is also designed to handle arbitrary spherical Earth density models.
In this paper, we describe how we implement the Earth in section II. Subsection III.1 discusses the techniques used to compute the eigenvalues and the eigenvectors, specifically what we call the internal eigenvectors. We then discuss how we take the internal amplitude constructed from those eigenvalues and eigenvectors and combine it into a probability in subsection III.2. This completes the calculation for atmospheric neutrinos. The solar and supernova calculations require several additional steps discussed in section IV. Numerical results for a variety of cases are shown in section V with additional numerical results shown in appendix A. We briefly describe the code usage in section VI; see also the more detailed user’s guide on the github page. In section VII we perform detailed speed and precision tests and make recommendations on how to optimally use the code depending on one’s usage case and we conclude in section VIII.
@@ II Earth
This paper is focused on the problem of neutrinos propagating through the Earth. Neutrinos propagating in other dense environments such as the Sun, supernova, the early Universe, and other celestial bodies each have their own unique challenges depending on particle content, density, typical neutrino energies, geometry, and temporal variation. For this discussion, the matter density of protons, neutrons, and electrons in the Earth are significant enough to appreciably modify neutrino oscillations, but not so large as to lead to significant absorption, which does not become relevant until $E\gtrsim\mathcal{O}($TeV)), at which point there are no significant oscillations in the standard three-flavor picture.
@@ II.1 Earth Models
The majority of the information we have about the interior of the Earth comes from seismic information. In the past, models of the Earth’s density profile assume spherical symmetry, but recently there has been some evidence of some small non-symmetric substructures [54, 55, 56, 57, 58, 59, 60, 61]. Nonetheless, we will focus on spherically symmetric Earth models as this leads to some computational advantages and the expected future sensitivity to such substructures in the Earth with atmospheric [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82], solar [83, 84], or supernova [85, 83, 86] neutrino oscillations is weak. A more complicated asymmetric Earth model would require a nontrivial modification of the provided code, but certainly is possible and many of the advantages of this algorithm would remain.
We consider the Earth’s density profile to be smoothly varying except at a finite number of sharp discontinuities. We base our density models on the Preliminary Reference Earth Model (PREM) [87] and use the electron fractions from [88]. We assume the atmosphere to be true vacuum.
@@ II.2 Trajectories
As the neutrinos pass from the production point, through the atmosphere, into the Earth, across the Earth and to the detector, the densities that they experience vary considerably. We break the trajectory up into four main segments. First, we start from the production point in the atmosphere to the surface of the Earth. As this section is purely in vacuum it is very easy to compute and modify. Second, we go from the surface of the Earth to the detector depth. Third, we go from the detector depth to the innermost point in the Earth. Finally, we go from the innermost point in the Earth to the detector depth at the actual detector which is the same as the previous step, up to a transpose for a spherically symmetric Earth. Reusing the calculations for one of these segments from the other saves a sizable amount of computational effort.
@@ III.1 One Layer Amplitude
$\sum_{i}V_{\alpha i}V^{*}_{\beta i}e^{-i\Delta\lambda_{i}L/(2E)}\,$, (2)
@@
using the shorthand $s_{ij}=\sin\theta_{ij}$ and $c_{ij}=\cos\theta_{ij}$. This choice is equivalent to the PDG choice [89], since if one right multiplies by $D_{3}(-\delta)$, one obtains the PDG PMNS matrix. As this is a change of the phase of $\nu_{3}$, no oscillation physics observable can be sensitive to this phase change. As will be shown shortly, we have a specific computational reason for making the above choice for the PMNS matrix.
@@
where $M^{2}=\text{Diag}(0,\Delta m^{2}_{21},\Delta m^{2}_{31})$ and the Wolfenstein matter potential is given by $W\equiv\text{Diag}(A_{mat},0,0)$ with $A_{mat}\equiv\pm 2\sqrt{2}\,G_{F}N_{e}E$ (+ for neutrinos and - for anti-neutrinos; the sign of $\delta$ also flips between neutrinos and antineutrinos). This Hamiltonian is a Hermitian matrix with complex entries because the $D_{3}(\delta)$ rotation. Since $D_{3}(\delta)$ commutes with matter potential $W$, and actually $R_{23}(\theta_{23})D_{3}(\delta)$ commutes with the matter potential, we separate out the $R_{23}(\theta_{23})D_{3}(\delta)$ rotation for the majority of the calculation, and then apply it at the very end of the full amplitude calculation. This is the maximal amount of information that can be rotated out of the Hamiltonian.
@@
which is a real symmetric matrix which not only has real eigenvalues, the $\lambda$’s, but the eigenvectors $V_{\alpha i}$ can also be chosen to be all real [^2] We have suppressed the arguments of the rotation matrices for readability.. We call this basis the “tilde” basis, see e.g. [38], which is related to the flavor basis by
@@
where $D_{3}^{*}(\delta)=D_{3}(-\delta)$. With this choice the layer oscillation amplitude, eq. 1 is a complex symmetric matrix, therefore only 6 of the 9 elements need to be explicitly calculated.
@@
which are all real and the relevant components of the normalized eigenvectors are calculated using [^3] The symbol “Adj[X]” stands for Adjugate of the matrix X, i.e. the transpose of the co-factor matrix of X.
@@
with $\Lambda\equiv\text{Diag}(\lambda_{1},\lambda_{2},\lambda_{3})$. Note that the eigenvalues of $H$ and $\widetilde{H}$ are the same since W commutes with $R_{23}D_{3}$ as noted earlier. Further details of this diagonalization method can be found in appendix B.
III.1.1 Calulation of Eigenvalues
@@
We provide two separate ways to find the solutions to this equation, one exact method by Cardano from 1545 and another that starts with an excellent approximation and then is improved by using Newton-Raphson (NR) iterations. With two iterations, this method has the same accuracy as the Cardano method using double precision arithmetic with a modest saving in computer resources. We begin by computing $\lambda_{3}$ either exactly or approximately and then using $\lambda_{3}$ to determine $\lambda_{1}$ and $\lambda_{2}$.
@@
The exact analytic expression for $\lambda_{3}$ is [22, 20, 21]
@@
with $n=0$ for NO and $n=1$ for IO. As one can see this requires two costly trigonometric function calls and one square root call.
@@
A simple but accurate approximate solution, given by Denton, Minakata and Parke (DMP), is [37, 38]
@@
$\text{with}~x$ $\equiv\frac{A_{mat}}{\Delta m^{2}_{ee}}~\text{and}~\Delta m^{2}_{ee}\equiv\Delta m^{2}_{31}-s^{2}_{12}\,\Delta m^{2}_{21}\,.$ (26)
$\Delta m^{2}_{ee}$ is the $\Delta m^{2}$ measured in $\nu_{e}$ disappearance experiments as defined in [42], see also [44]. This approximation for $\lambda_{3}$ is an excellent one for both mass orderings with a fractional difference to the exact eigenvalue of better than $10^{-4}$ for all neutrino energies, and is exact in vacuum. Also it contains just one square root and no trigonometric functions.
The precision of eq. 26 can be further improved to arbitrary precision by using one or more NR iterations [18] [^4] Other improvement methods could be used here, such as second order NR, however we use 1st order NR because of its simplicity and rapid rate of convergence.:
@@
where $\chi$ is the characteristic equation (eq. 23) and $\chi^{\prime}=3\lambda^{2}-2A\lambda+B$ is the derivative with respect to $\lambda$.
@@ III.2 Amplitudes to Probabilities
Given the inner amplitude in the tilde basis (without the $\theta_{23}$ and $\delta$ rotation) for one layer, we next combine the amplitudes through each layer in the Earth. We also consider that the trajectories may start in the atmosphere and that the detector may be inside the Earth’s crust.
@@ III.2.1 Combining the Layer Amplitudes
Each of these segments is then sub-divided into multiple layers with different densities and different path lengths in each layer depending on the zenith angle. This construction is designed in this way such that the oscillation amplitudes in the $I$ and the $O$ segments are the transpose of one another since the density-length profiles are identical except in inverse order which can save up to a factor of two in computational effort. This construction also factorizes out the propagation in the atmosphere which can be trivially calculated and adjusted for different production heights.
@@
There are a total of $m+n+1$ multiplications of $3\times 3$ complex matrices, down from $m+2n+1$ if the identity $\widetilde{A}_{O}=\widetilde{A}_{I}^{T}$ was not used. When the detector is within a few kilometers of the Earth’s surface, $m$ is expected to be a small integer and $n$ a much larger integer, especially for trajectories that cross through much of the Earth. If we are using an N-layer Earth model, then for any trajectory $m+n\leq N$, with the maximum number of such matrix multiplications occurring for a core crossing trajectory.
At this point $\theta_{23}$ and $\delta$ have not appeared in the calculation of $\widetilde{\cal A}_{traj}$, which is the most computationally expensive part of computing the entire probability. Thus we can now calculate the oscillation probability with many values of $\sin^{2}\theta_{23}$ and $\delta$ very inexpensively. This is especially useful as these variables are the variables which are known with the least precision at the current time.
@@ III.2.2 Transforming back to the flavor basis
To include $\theta_{23}$ and $\delta$ and transform back to the flavor basis, we perform the following rotation on the amplitude along the full trajectory through the Earth $\widetilde{\cal A}_{traj}$:
$\widetilde{{\cal A}}_{f}=R_{23}D_{3}\widetilde{\cal A}_{traj}D_{3}^{*}R_{23}^{T}\,$, (36)
@@
$P_{\alpha\beta}=|\widetilde{{\cal A}}_{f}[\alpha,\beta]|^{2}\,.$ (37)
It is straightforward to confirm that the sum of each row adds to 1 and the sum of each column adds to 1, due to the unitarity of $\widetilde{{\cal A}}_{f}$.
However, due to this unitary nature of the P matrix, only 4 elements need to be directly calculated from $\tilde{\cal A}_{traj}$, saving on complex computation. We have chosen the simplest independent four;
@@
Eqs. 38-41 are identical to those found in ref. [90, 91, 92]. Since $\widetilde{{\cal A}}_{traj}$ is unitary, all oscillation probabilities can be calculated using only 4 independent elements of $\widetilde{{\cal A}}_{traj}$. Here, our choices are determined so as to minimize complex arithmetic.
@@
$P_{\tau\tau}$ $=P_{ee}+P_{\mu\mu}+P_{e\mu}+P_{\mu e}-1\,.$
A basic cross-check is that all elements of P must satisfy, $0\leq P_{\alpha\beta}\leq 1$. When $\widetilde{{\cal A}}_{traj}$ is unitary, this set of constraints is guaranteed to be satisfied.

### v2

@@ I Introduction
To address this issue, for long-baseline accelerator and reactor neutrino oscillations, we developed the NuFast-LBL algorithm [18] and the associated code on github [19]. This algorithm was based on a variety of theoretical neutrino oscillation results obtained over a number of years including some of the earliest full expressions for neutrino oscillations in constant matter [20, 21], solutions to the cubic equation [22, 18], the role of CP violation [23, 24, 25, 26, 27, 28, 29], the optimal means of computing the eigenvectors [30, 31, 32, 33, 34, 18, 35], and the optimal structure for the oscillation probabilities as well as various other theoretical advances [36, 37, 38, 39, 40], see also [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
In this paper, we use parts of the NuFast-LBL algorithm and modify it to handle the varying density profiles through the Earth in a novel calculation. The additional complexities of traversing sharply varying density profiles require a new approach constructed from the ground up, which is the result presented in this paper. We provide a new code called NuFast-Earth, also available on github, see [54] [^1] The code is available at github.com/PeterDenton/NuFast-Earth., which is significantly more powerful than the long-baseline (LBL) code. Notably, it intelligently saves reused aspects of the calculations for future uses which vastly speeds up many calculations of the oscillation probabilities. It is also designed to handle arbitrary spherical Earth density models with numerous models coded up and the ability to easily add more.
In this paper, we describe how we implement the Earth in section II. Subsection III.1 discusses the techniques used to compute the eigenvalues and the eigenvectors, specifically what we call the internal eigenvectors, with calculational details given in appendix A. We then discuss how we take the internal amplitude constructed from those eigenvalues and eigenvectors and combine it into a probability in subsection III.2. This completes the calculation for atmospheric neutrinos. The solar and supernova calculations require several additional steps discussed in section IV. Numerical results for a variety of cases are shown in section V with additional numerical results shown in appendix B. We briefly describe the code usage in section VI; see also the more detailed user’s guide on the github page. In section VII we perform detailed speed and precision tests and make recommendations on how to optimally use the code depending on one’s usage case and we conclude in section VIII.
@@ II Earth
This paper is focused on the problem of neutrinos propagating through the Earth. Neutrinos propagating in other dense environments such as the Sun, supernova, the early Universe, and other celestial bodies each have their own unique challenges depending on particle content, density, typical neutrino energies, geometry, and temporal variation. For this discussion, the matter density of protons, neutrons, and electrons in the Earth are significant enough to appreciably modify neutrino oscillations, but not so large as to lead to absorption, which does not become relevant until $E\gtrsim\mathcal{O}($TeV)), at which point there are oscillations are extremely damped in the standard three-flavor picture.
@@ II.1 Earth Models
The majority of the information we have about the interior of the Earth comes from seismic information. In the past, models of the Earth’s density profile assume spherical symmetry, but recently there has been some evidence of some small non-symmetric substructures [55, 56, 57, 58, 59, 60, 61, 62]. Nonetheless, we will focus on spherically symmetric Earth models as this leads to some computational advantages and the expected future sensitivity to such substructures in the Earth with atmospheric [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83], solar [84, 85], or supernova [86, 84, 87] neutrino oscillations is weak. A more complicated asymmetric Earth model could be implemented in the code via the single trajectory mode, but is not the focus of this paper.
We consider the Earth’s density profile to be smoothly varying except at a finite number of sharp discontinuities. We base our density models on the Preliminary Reference Earth Model (PREM) [88] and use the electron fractions from [89]. We assume the atmosphere to be true vacuum [^2] This is an excellent approximation. We have calculated the maximum shift in the probabilities across $E\in[2,40]$ GeV and $\cos\theta_{z}\in[-1,0.1]$ to be: $\max|\Delta P|$ $\nu_{\mu}\to\nu_{\mu}$ $\nu_{\mu}\to\nu_{e}$ $\nu$ $\bar{\nu}$ $\nu$ $\bar{\nu}$ NO $5\times 10^{-8}$ $3\times 10^{-8}$ $3\times 10^{-7}$ $3\times 10^{-7}$ IO $8\times 10^{-8}$ $4\times 10^{-8}$ $3\times 10^{-7}$ $3\times 10^{-7}$ where the production height is 10 km, the detector depth is 2 km, and the atmosphere is either true vacuum or has $\rho Y_{e}=0.5\times 10^{-3}$ g/cm3 which is conservative, e.g. large..
@@ II.2 Trajectories
As the neutrinos pass from the production point, through the atmosphere, into the Earth, across the Earth and to the detector, the densities that they experience vary considerably. We break the trajectory up into four main segments. First, we start from the production point in the atmosphere to the surface of the Earth. As this section is purely in vacuum it is very easy to compute and modify. Second, we go from the surface of the Earth to the detector depth. Third, we go from the detector depth to the innermost point in the Earth. Finally, we go from the innermost point in the Earth to the detector depth at the actual detector which is the same as the previous step, up to a transpose for a spherically symmetric Earth. Reusing the calculations for one of these segments from the other saves a sizable amount of computational effort. These trajectories are shown in Fig. 1 and are discussed in more detail below.
@@ III.1 One Layer Amplitude
$\sum_{i}V_{\alpha i}V^{*}_{\beta i}e^{-i\lambda_{i}L/(2E)}\,$, (2)
@@
using the shorthand $s_{ij}=\sin\theta_{ij}$ and $c_{ij}=\cos\theta_{ij}$. This choice is equivalent to the PDG choice [90], since if one right multiplies by $D_{3}(-\delta)$, one obtains the PDG PMNS matrix, see, e.g., [91]. As this is a change of the phase of $\nu_{3}$, no oscillation physics observable can be sensitive to this phase change. As will be shown shortly, we have a specific computational reason for making the above choice for the PMNS matrix.
@@
where $M^{2}=\text{Diag}(0,\Delta m^{2}_{21},\Delta m^{2}_{31})$ and the Wolfenstein matter potential is given by $W\equiv\text{Diag}(A_{mat},0,0)$ with $A_{mat}\equiv\pm 2\sqrt{2}\,G_{F}N_{e}E$ (+ for neutrinos and - for antineutrinos; the sign of $\delta$ also flips between neutrinos and antineutrinos). This Hamiltonian is a Hermitian matrix with complex entries because the $D_{3}(\delta)$ rotation. Since $R_{23}(\theta_{23})D_{3}(\delta)$ commutes with the matter potential, we separate out the $R_{23}(\theta_{23})D_{3}(\delta)$ rotation for the majority of the calculation, and then apply it at the very end of the full amplitude calculation. This is the maximal amount of information that can be rotated out of the Hamiltonian.
@@
which is a real symmetric matrix that not only has real eigenvalues, the $\lambda$’s, but the eigenvectors $\widetilde{V}_{\alpha i}$ can also be chosen to be all real [^3] We have suppressed the arguments of the rotation matrices for readability.. We call this basis the “tilde” basis, see e.g. [39], which is related to the flavor basis by
@@
where $D_{3}^{*}(\delta)=D_{3}(-\delta)$. With this choice the oscillation amplitude in one layer, eq. 1, is a complex symmetric matrix, therefore only 6 of the 9 elements need to be explicitly calculated.
@@
which are all real and the relevant components of the normalized eigenvectors are calculated using [^4] The symbol “Adj[X]” stands for Adjugate of the matrix X, i.e. the transpose of the cofactor matrix of X.
@@
with $\Lambda\equiv\text{Diag}(\lambda_{1},\lambda_{2},\lambda_{3})$. Note that the eigenvalues of $H$ and $\widetilde{H}$ are the same since W commutes with $R_{23}D_{3}$ as noted earlier. Further details of this diagonalization method can be found in appendix A.
III.1.1 Calculation of Eigenvalues
@@
We provide two separate ways to find the solutions to this equation, one exact method by Cardano from 1545 and another that starts with an excellent approximation and then is improved to arbitrary precision by using Newton-Raphson (NR) iterations. With two iterations, this method has the same accuracy as the Cardano method using double precision arithmetic with a modest saving in computer resources. We begin by computing $\lambda_{3}$ either exactly or approximately and then using $\lambda_{3}$ to determine $\lambda_{1}$ and $\lambda_{2}$.
@@
The exact analytic expression for $\lambda_{3}$ is
@@
with $n=0$ for NO and $n=1$ for IO [22, 20, 21]. As one can see this requires two costly trigonometric function calls and one square root call.
@@
A simple but accurate approximate solution is
@@
$\text{with}~x$ $\equiv\frac{A_{mat}}{\Delta m^{2}_{ee}}~\text{and}~\Delta m^{2}_{ee}\equiv\Delta m^{2}_{31}-s^{2}_{12}\,\Delta m^{2}_{21}\,$, (26)
from [38, 39]. $\Delta m^{2}_{ee}$ is the $\Delta m^{2}$ measured in $\nu_{e}$ disappearance experiments as defined in [43], see also [45]. This approximation for $\lambda_{3}$ is an excellent one for both mass orderings with a fractional difference to the exact eigenvalue of better than $10^{-4}$ for all neutrino energies, and is exact in vacuum. Also it contains just one square root and no trigonometric functions.
The precision of eq. 26 can be further improved to arbitrary precision by using one or more NR iterations [^5] Other improvement methods could be used here, such as second order NR, however we use 1st order NR because of its simplicity and rapid rate of convergence.:
@@
where $\chi$ is the characteristic equation (eq. 23) and $\chi^{\prime}=3\lambda^{2}-2A\lambda+B$ is the derivative with respect to $\lambda$ [18].
@@ III.2 Amplitudes to Probabilities
Given the inner amplitude in the tilde basis (without the $\theta_{23}$ and $\delta$ rotation) for one layer, we next combine the amplitudes through each layer in the Earth. We also consider that the trajectories may start in the atmosphere and that the detector may be located inside the Earth.
@@ III.2.1 Combining the Layer Amplitudes
Each of these segments is then subdivided into multiple layers with different densities and different path lengths in each layer depending on the zenith angle. This construction is designed in this way such that the oscillation amplitudes in the $I$ and the $O$ segments are the transpose of one another since the density-length profiles are identical except in reverse order which can save up to a factor of two in computational effort. This construction also factorizes out the propagation in the atmosphere which can be trivially calculated and adjusted for different production heights.
@@
There are a total of $m+n+1$ multiplications of $3\times 3$ complex matrices, down from $m+2n+1$ if the identity $\widetilde{A}_{O}=\widetilde{A}_{I}^{T}$ was not used. When the detector is within a few kilometers of the Earth’s surface, $m$ is expected to be a small integer and $n$ a much larger integer, especially for trajectories that cross through much of the Earth, leading to nearly a factor of two reduction in computational cost. If we are using an $N$-layer Earth model, then for any trajectory $m+n\leq N$, with the maximum number of such matrix multiplications occurring for a core crossing trajectory.
At this point $\theta_{23}$ and $\delta$ have not appeared in the calculation of $\widetilde{\cal A}_{traj}$, which is the most computationally expensive part of computing the entire probability. Thus we can now calculate the oscillation probability with many values of $\sin^{2}\theta_{23}$ and $\delta$ very inexpensively. This is especially useful as these variables are the oscillation parameters which are known the least at the current time, and both play an important role in atmospheric oscillations.
@@ III.2.2 Transforming back to the flavor basis
To include $\theta_{23}$ and $\delta$ and transform back to the flavor basis, we perform the following rotation on the amplitude along the full trajectory through the Earth and the atmosphere $\widetilde{\cal A}_{traj}$:
${\cal A}_{f}=R_{23}D_{3}\widetilde{\cal A}_{traj}D_{3}^{*}R_{23}^{T}\,$, (36)
@@
$P_{\alpha\beta}=|\,{{\cal A}}_{f}[\alpha,\beta]\,|^{2}\,.$ (37)
It is straightforward to confirm that the sum of each row adds to 1 and the sum of each column adds to 1, due to the unitarity of ${{\cal A}}_{f}$.
Due to this unitary nature of the P matrix, only 4 elements need to be directly calculated from $\tilde{\cal A}_{traj}$, saving on complex computation. We have determined the simplest independent four;
@@
Eqs. 38-41 are identical to those found in refs. [92, 93, 94]. Since $\widetilde{{\cal A}}_{traj}$ is unitary, all oscillation probabilities can be calculated using only 4 independent elements of $\widetilde{{\cal A}}_{traj}$. Here, our choices are determined so as to minimize complex arithmetic.
@@
$P_{\tau\tau}$ $=P_{ee}+P_{\mu\mu}+P_{e\mu}+P_{\mu e}-1\,$,
completing the calculation. A basic cross-check is that all elements of $P$ must satisfy, $0\leq P_{\alpha\beta}\leq 1$. When $\widetilde{{\cal A}}_{traj}$ is unitary, this set of constraints is guaranteed to be satisfied.